OpenStudy (anonymous):

(-2)^4+3 square root 120-20/4^3+|5(-1)|

OpenStudy (lasttccasey):

Is this correct?\[(-2)^4+3 \sqrt{120}-20/4^3+|5×(-1)|\]

OpenStudy (anonymous):

Yes

OpenStudy (lasttccasey):

Okay, are you familiar with PEMDAS?

OpenStudy (anonymous):

No

OpenStudy (lasttccasey):

PEMDAS is an acronym for the order of operations. That is the order we are supposed to work math problems such as this. P=Parenthesis E=Exponential M & D = Multiplication & Division A & S = Addition & Subtraction the M&D + A&S are to be done in order from left to right, whichever comes first.

OpenStudy (anonymous):

Okay

OpenStudy (lasttccasey):

First step is do all parenthesis: \[(-2)^4+3 \sqrt{120}-20/4^3+abs(5×(-1))\]\[16+3 \sqrt{120}-20/4^3+5\]

OpenStudy (lasttccasey):

Next Exponentials: \[16+3 \sqrt{120}-20/64+5\]

OpenStudy (anonymous):

Where did 64 come from?

OpenStudy (lasttccasey):

the 4^3

OpenStudy (anonymous):

Ok

OpenStudy (lasttccasey):

Alright then just multiple and divide from left to right. I'm also going to simplify the radical but instead of dividing 20/64 I'm just going to simplify to avoid decimals. \[16+6 \sqrt{30}-5/16+5\]

OpenStudy (lasttccasey):

Final step is to add and subtract. \[6\sqrt{30}+331/16\]

OpenStudy (lasttccasey):

decimal form answer: 53.55085

OpenStudy (anonymous):

Thank you

OpenStudy (lasttccasey):

You're welcome.